OpenStudy (anonymous):

Solve x + 3y = 9 3x – 3y = –13 A. (–1, ten thirds) B. (1, negative ten thirds) C. (–1, 3 over 10) D. (1, negative 3 over 10)

OpenStudy (primeralph):

Add both equations.

OpenStudy (anonymous):

answer is A?

OpenStudy (anonymous):

the answer is (A). if you add the two expressions: x + 3x = 4x 3y + (-3y) = 0 9 + (-13) = -4 you end up with: 4x = -4, so x = -1 You plug -1 into either one of the original expression and you should end up with y = 10/3

OpenStudy (anonymous):

Is this algebra 1

OpenStudy (primeralph):

@juliaann You might want to learn to do these now; we wont always be there to give you the answers.

OpenStudy (anonymous):

algebra 2! @Courage.

OpenStudy (anonymous):

Summer class

OpenStudy (anonymous):

A

OpenStudy (anonymous):

@primeralph yes i know I only post the ones i have a hard time with